Solve by factoring.
x squared - 9x - 4 = 6

4x squared - 8x + 3 = 5

Answer by stanbon(75887)

Solve by factoring.
x squared - 9x - 4 = 6
x^2 - 9x -10 = 0
(x-10)(x+1) = 0
x = 10 or x = -1
----------------------
4x squared - 8x + 3 = 5
4x^2 -8x -2 = 0
2x^2-4x-1 = 0
x = [4 +- sqrt(16-4*2*-1)]/4
x = [4 +- sqrt(24)]/4
x = [4 +- 2sqrt(6)]/4
x = 1+(1/2)sqrt(6) or x = 1-(1/2)sqrt(6)
=============================
Cheers,
Stan H.